A firm operates in a perfectly competitive market. The market price of its product is 4 birr and the total cost function is given by TC= 1/3Q3 -5Q2+20Q+50, where TC is the total cost and Q is the level of output.
1
Expert's answer
2020-10-19T13:39:52-0400

Solution

Cost function

"Tc=\frac{1}{3}q^3-5q^2+20q+50"

At maximize level of price

MC=MCP

"MC=\frac{d(Tc)}{dq}"

"MC=\frac{d}{dq}(\frac{1}{3}q^3-5q^2+20q+50)"


"MC=q^2-10q+20"


And MCP=4

So

"q^2-10q+20=4\\q^2-10q+16=0"

So

q=8 and 2

Maximize level of price will be 2 because "\frac{d^2(Tc)}{dq}<0" for 2.